OpenStudy (anonymous):

Will medal!! A student 98 and 87 on his first two quizzes. Use a compound inequality to find the possible value for the third quize that would give him a average between 90 and 95. A.87=

OpenStudy (anonymous):

The question can be written as \[90 \leq \frac{98+87+x}3 \leq 95\] which means \[\frac{98+87+x_{low}}3 = 90\] \[\frac{98+87+x_{high}}3 = 95\] if we define \(x_{low}\) and \( x_{high}\) by how the numbers in the answer \[x_{low} \leq x \leq x_{high}\]

OpenStudy (anonymous):

So B...

OpenStudy (anonymous):

No

OpenStudy (anonymous):

If you choose B \[\frac{98+87+95}3 = 93\frac13\] \[\frac{98+87+90}3 = 91\frac23\] The student will average between 91.6666 and 93.3333, but the question wants the student to average between 90 and 95

OpenStudy (anonymous):

So he would have to get higher than 95 on at least one test?

OpenStudy (anonymous):

His average on the first two tests is \[\frac{98+87}2 = 92.5\] So to raise the average to 95, the third test score would have to be above 95 To lower the average to 90, the third test score would have to be below 90 You can work out the two limits of test scores by solving for \(x_{low}\) and \(x_{high}\) in my first reply

OpenStudy (anonymous):

So the solution is 87=<x=<98

OpenStudy (anonymous):

No \[\frac{98+87+98}3 = 94.3333...\] \[\frac{98+87+87}3 = 90.6666...\]

OpenStudy (anonymous):

Why not just solve the equations? \[\frac{98+87+x_{low}}3 = 90\] \[98+87+x_{low} = 3\times90\] \[x_{low} = 3\times90 - 98 - 87\] similarly \[x_{high} = 3\times95 - 98 - 87\]

OpenStudy (anonymous):

I don't know how to solve using xhigh and xlow

OpenStudy (anonymous):

They're just numbers. They're nothing special. I wrote them because the answer defines the third test score to be x, and you want to find the higher and lower limits of x. I could write \[x_{low} = \text{lower limit of third test score}\] \[x_{high} = \text{higher limit of third test score}\] so the question would be \[\frac{98+87+\text{lower limit of third test score}}3 = 90\] \[\frac{98+87+\text{higher limit of third test score}}3 = 95\]

OpenStudy (anonymous):

I working on solving, So far I have xhigg=100. Is that correct.

OpenStudy (anonymous):

\[\frac{98+87+100}3 = 95\] Yes

OpenStudy (anonymous):

So xlow=85 and xhigh=100. Correct?

OpenStudy (anonymous):

\[\frac{98+87+85}3 = 90\] Yeah, that's right :)

OpenStudy (anonymous):

THANK YOU SO MUCH!!!! This is a real life saver!!